Modern Physics Tunnel Effect HARD problem

A beam of electrons is incident on a barrier 6 eV high and.2nm wide.Find the energy they should have if 1.00percent of them are to get through thebarrier.The equations that have to be used arePlease help
Answer #1
The transmission coefficient is 1% = 0.01
0.01 = e-2k2 L
- 2 K2 L =ln ( 0.01 )
- 2 K2L= -4.6051
K2 = 4.6051/ 2 (0.2 x10- 9)
=1.1513x 1010
√[2m(U -E ) = h(bar)1.1513x 1010
[2m(U -E )= [h(bar)1.1513x 1010]2
(U -E )= [h(bar)1.1513x1010]2/ 2m
here m = 9.1 x 10-31 kg, h (bar) = 1.0536 x10-34
U - E = 8.0756x10-19J
In eV
U - E=8.0756x10-19J( 1eV/ 1.6x10-19 J )
= 5.0473eV
E = U - 5.0473eV
= 6eV -5.0473eV
= 0.9526eV